# Minkowski tensors for voxelized data: robust, asymptotically unbiased estimators

### v1

@@ Abstract
Minkowski tensors, also known as tensor valuations, provide robust $n$-point information for a wide range of random spatial structures. Local estimators for voxelized data, however, are unavoidably biased even in the limit of infinitely high resolution. Here, we substantially improve a recently proposed, asymptotically unbiased algorithm to estimate Minkowski tensors for voxelized data. Our improved algorithm is more robust and efficient. Moreover we generalize the theoretical foundations for an asymptotically bias-free estimation of the interfacial tensors to the case of finite unions of compact sets with positive reach, which is relevant for many applications like rough surfaces or composite materials. As a realistic test case, we consider, among others, random (beta) polytopes. We first derive explicit expressions of the expected Minkowski tensors, which we then compare to our simulation results. We obtain precise estimates with relative errors of a few percent for practically relevant resolutions. Finally, we apply our methods to real data of metallic grains and nanorough surfaces, and we provide an open-source python package, which works in any dimension.
@@ 1 Introduction
In all of these cases, Minkowski tensors (also known as tensor valuations) [55, 28] provide a comprehensive, unified shape analysis. The Minkowski tensors [57] are generalizations of the well-known Minkowski functionals (or intrinsic volumes) from integral geometry [55]; see Section 2 for more details and the theoretical background. The intrinsic volumes represent intuitive geometric information, e.g., on the volume, surface area, or curvature; the Minkowski tensors additionally distinguish different orientations with respect to these geometrical properties [37].
@@
A challenge, however, is the application of Minkowski tensors to voxelized data because the discrete set of available orientations introduces a systematic bias usually even in the limit of infinitely high resolution. For gray-scale data, marching-cube algorithms can turn the voxelized data into triangulated meshes with continuous orientations, which results in asymptotically bias-free algorithms [39, 38, 61, 62]. If no gray-scale data is available, any local algorithm on black-and-white voxels will be asymptotically biased, i.e., the estimator will not converge to the true value in the limit of infinite resolution [63].
@@
An algorithmic solution to the problem was recently proposed in [21], which had been inspired by [43]. The key idea is to construct Voronoi cells that can encode global information and thus guarantee asymptotically bias-free estimators of all Minkowski tensors for sets of positive reach. This requirement of a positive reach, however, excludes many interesting applications, like rough surfaces and composite materials. More importantly, the proposed algorithm suffered from numerical instabilities since it relies on a matrix inversion.
@@
Here, we substantially advance the robustness of the algorithm by replacing the matrix inversion by a least squares fit. We, moreover, increase its efficiency by using an unbiased estimator of the Voronoi volume integrals. This revised algorithm is also asymptotically unbiased for sets of positive reach, as we prove rigorously. We demonstrate its robustness and accuracy for exemplary cases of geometric shapes, stochastic models, and real data.
@@
Even for a simple example of a set without positive reach, our algorithm achieves surprisingly good results (with relative deviations of less than 3% in the eigenvalues of the interfacial tensors, see Table 6). Since such precision is not theoretically substantiated, we also introduce an alternative method to estimate interfacial tensors, for which we are able to prove that it is asymptotically unbiased even for finite unions of compact sets with positive reach.
@@
Together with the paper, we publish an open-source python package including both methods. It is the first implementation of Minkowski tensors that we know of that can be applied to any dimension. It can also be applied to any discrete representation of the set, i.e., not only for pixelated images, but representations using non-cubic lattice or even random point patterns. The last generalization is especially helpful in high dimensions, where similar to numerical integration, lattice representations may converge too slowly. Details on the algorithm, implementation, and suitable choices of parameters are discussed in Sections 4.14.3.
@@
We demonstrate the reliability of our methods for simple geometric test cases in Sections 4.4 and 4.5. For a stochastic test case, we apply our estimator to a classic example of random convex sets, isotropic random polytopes (see Section 5), both in theory (see Section 5.1) and simulations (see Section 5.2).
@@
In Section 6, we turn to experimental data. First, we analyze metallic grains in a polycrystalline nickel-based superalloy using data from [60]. The Minkowski tensors characterize the cells to be distinctly more anisotropic with respect to curvature than the surface area (see Section 6.1). Then, we determine the interfacial tensors of nanorough surfaces from [59] (see Section 6.2).
@@ 2 Minkowski tensors in integral geometry
[…]
  • They form a basis of the vector space of continuous, isometry invariant, additive, functions on ${\mathcal{K}}^{d}$; $V_{i}$ is positively homogeneous of degree $i$ [26, Theorem 4.20].
  • They arise as coefficient functionals of a Steiner formula [26, Theorem 3.10].
  • They satisfy various integral geometric formulas such as Crofton formulas, projection formulas and (intersectional as well as additive) kinematic formulas [26, Chap. 5].
  • They can be additively extended to polyconvex sets (finite unions of compact, convex sets) and play a key role in the classical theory of geometric valuations (see [54, Chap. 1] and [26, Chap. 4.5]).
The intrinsic volumes have also been called Minkowski functionals (or quermassintegrals, though with a different normalization and notation). Due to their properties, the Minkowski functionals are useful descriptors that can be applied for the analysis of complex spatial structure.
Since the Minkowski functionals are isometry invariant, their usefulness in characterizing anisotropic features of objects under investigation is limited. The Minkowski tensors (see [25, Chap. 2]), which we introduce next, provide a more general set of tensor-valued geometric descriptors that are sensitive to position and orientation of geometric objects in Euclidean space. We start by providing a brief introduction to symmetric tensors and tensor-valued functionals as needed for the present purpose.
@@
where $m_{k}:=|\{\ell\in[r]:i_{\ell}=k\}|\in\{0,\ldots,r\}$, $k=1,\ldots,d$, counts how often the number $k$ appears among the indices $i_{1},\dots,i_{r}$. In particular, for $d=r=2$, we have $T(e_{1},e_{1})=t_{11}$, $T(e_{2},e_{2})=t_{22}$ and $T(e_{1},e_{2})=T(e_{2},e_{1})=\frac{1}{2}t_{12}$, hence $T=T(e_{1},e_{1})e_{1}^{2}+2T(e_{1},e_{2})e_{1}e_{2}+T(e_{2},e_{2})e_{2}^{2}$ with $2T(e_{1},e_{2})e_{1}e_{2}=T(e_{1},e_{2})e_{1}e_{2}+T(e_{2},e_{1})e_{2}e_{1}$.
@@
In order to introduce the basic Minkowski tensors of (nonempty) convex bodies, we use the support measures which can be considered as local versions of the intrinsic volumes. For a convex body $K\subset{\mathbb{R}}^{d}$, the support measure $\Lambda_{j}(K,\cdot)$ is a Borel measure on ${\mathbb{R}}^{d}\times\mathbb{S}^{d-1}$ which satisfies a local Steiner formula and $\Lambda_{j}(K,{\mathbb{R}}^{d}\times\mathbb{S}^{d-1})=V_{j}(K)$ for $j\in\{0,\ldots,d-1\}$. For basic properties of the support measures and a representation in the case where $K$ is a polytope, we refer to [53, Chapter 4.2]. The maps $K\mapsto\Lambda_{j}(K,\cdot)$ are measure-valued, additive (valuations), weakly continuous and for Borel sets $\alpha\subseteq{\mathbb{R}}^{d}$ and $\beta\subseteq\mathbb{S}^{d-1}$ they satisfy the covariance condition $\Lambda_{j}(gK,g\alpha\times g_{0}\beta)=\Lambda_{j}(K,\alpha\times\beta)$, where $g:{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$ is a rigid motion and $g_{0}$ is its rotational part (that is, $g(x)=g_{0}(x)+t$ for $x\in{\mathbb{R}}^{d}$, where $t\in{\mathbb{R}}^{d}$ is a translation vector and $g_{0}$ is a rotation). More generally, the support measures have been defined (as signed measures) for the more general class of sets of positive reach (see [49, Chapter 4] for an introduction) and for finite unions of compact sets with positive reach all of whose finite intersections have again positive reach (see [49, Chapter 5] and the literature cited there); the corresponding class of sets in ${\mathbb{R}}^{d}$ is denoted by $\mathcal{U}^{d}$. In fact, extensions of these measures to arbitrary closed subsets of ${\mathbb{R}}^{d}$ have been studied in [23, 24] and will be considered in Section 3.
If $K$ is a convex body or a compact set with positive reach, then the basic Minkowski tensors of $K$ are defined by
@@
$\Phi_{k}^{r,s}(K):=\frac{1}{r!s!}\frac{\omega_{d-k}}{\omega_{d-k+s}}\int_{{\mathbb{R}}^{d}\times\mathbb{S}^{d-1}}x^{r}u^{s}\,\Lambda_{k}(K,\mathrm{d}(x,u))$ (2.4)
for $k=0,\dots,d-1$ and $r,s\in{\mathbb{N}}_{0}$,
@@
Minkowski tensors can be defined not only for convex bodies or compact sets with positive reach. By additivity their domain can be extended to the class of polyconvex sets (finite unions of convex bodies). Moreover they can be defined also for sets from the class $\mathcal{U}^{d}$, since this generalization is also true for the support measures. Furthermore, Minkowski tensors can be introduced for even more general classes of sets provided that the support measures are defined in a natural way and the integrals exist (see Section 3 and the references given there). An explicit description of the Minkowski tensors of polytopes is recalled in Section 4.4, in Section 4.5 we evaluate formulas for nonconvex sets. In the case of sets with positive reach or for even more general classes of compact sets $K$ (such as arbitrary finite unions of sets with positive reach), the support measures $\Lambda_{j}(K,\cdot)$ are consistently replaced by the reach measures $\mu_{j}(K;\cdot)$ in definition (2.4); the relationship between these measures will be discussed in Section 3.
It is a straightforward consequence of the properties of the support measures of convex bodies that the Minkowski tensors of convex bodies are continuous, isometry (rigid motion) covariant (for a detailed definition, especially of the underlying notion of polynomial behavior with respect to translations, we refer to [25, Section 2.2]) and additive tensor functionals (valuations) on the space of convex bodies in ${\mathbb{R}}^{d}$. By a fundamental result due to Alesker, the vector space $T_{p}$ of all mappings $\Gamma:\mathcal{K}^{d}\to\mathbb{T}^{p}$ having these properties is spanned by the tensor valuations $Q^{m}\Phi_{k}^{r,s}$, where $m,r,s\in{\mathbb{N}}_{0}$ satisfy $2m+r+s=p$, where $k\in\{0,\dots,d\}$, and where $s=0$ if $k=d$. In contrast to the real-valued and vector-valued case, these tensor valuations are no longer linearly independent. A study of linear dependencies was initiated by McMullen and completed by Hug, Schneider, Schuster (for precise references and detailed statements, see, e.g., [25, Thms. 2.6 and 2.7]). These general investigations imply that if $T_{2,k}$ denotes the subspace of all maps in $T_{2}$ that are homogeneous of degree $k$, then $T_{2}=T_{2,0}\oplus\cdots\oplus T_{2,d+2}$ (direct sum decomposition), $\dim T_{2}=3d+1$, and a basis of $T_{2,k}$ is displayed in Table 1. Denoting by $T_{2}^{\ast}$ the subspace of translation invariant maps in $T_{2}$ and by $T_{2,k}^{\ast}$ the subspace of translation invariant maps in $T_{2,k}$, we have $T_{2}^{\ast}=T_{2,0}^{\ast}\oplus\cdots\oplus T_{2,d}^{\ast}$, $\dim T_{2}^{\ast}=2d$, and a basis of $T_{2,k}^{\ast}$ is displayed in Table 1.

### v2

# Minkowski tensors for point clouds and voxelized data: robust, asymptotically unbiased estimators

@@ Abstract
Minkowski tensors, also known as tensor valuations, provide robust $n$-point information for a wide range of random spatial structures. Local estimators for point clouds, e.g., representing voxelized data, however, are unavoidably biased even in the limit of infinitely high resolution. Here, we substantially improve a recently proposed, asymptotically unbiased algorithm to estimate Minkowski tensors from point clouds. Our improved algorithm is more robust and efficient. Moreover we generalize the theoretical foundations for an asymptotically bias-free estimation of the interfacial tensors, among others, to the case of finite unions of compact sets with positive reach, which is relevant for many applications like rough surfaces or composite materials. As a realistic test case of random spatial structures, we consider random (beta) polytopes. We first derive explicit expressions of the expected Minkowski tensors, which we then compare to our simulation results. We obtain precise estimates with relative errors of a few percent for practically relevant resolutions. Finally, we apply our methods to real data of metallic grains and nanorough surfaces, and we provide an open-source python package, which works in any dimension.
@@ 1 Introduction
In all of these cases, Minkowski tensors (also known as tensor valuations) [67, 36] provide a comprehensive, unified shape analysis. The Minkowski tensors [70] are generalizations of the well-known Minkowski functionals (or intrinsic volumes) from integral geometry [67]; see Section 2 for more details and the theoretical background. The intrinsic volumes represent intuitive geometric information, e.g., on the volume, surface area, or curvature; the Minkowski tensors additionally distinguish different orientations with respect to these geometrical properties [44]; see also Fig. 1.
Figure 1: Distinct Minkowski tensors quantify different types of anisotropy: (a–b) moment tensors of the volume and boundary distributions capture the anisotropy of disk arrangements; (c) the Euler characteristic $\chi$ is a topological constant, generalizations to surface tensors $\Phi_{0}^{0,2s}$ are proportional to $\chi$ and the metric tensor and hence always isotropic, see (2.8); (d–f) anisotropy of the surface‑normal distribution is measured by surface tensors, different symmetries are captured by tensors of a corresponding rank; (g) curvature anisotropy is incorporated by an additional curvature function to the surface tensors.
@@
A challenge, however, is the application of Minkowski tensors to point clouds, e.g., representing voxelized data, where a systematic bias can often not be avoided even in the limit of infinitely high resolution. In contrast, for voxelized gray-scale data, marching-cube algorithms can create triangulated meshes with continuous orientations, which results in asymptotically bias-free algorithms [51, 50, 75, 76]. If no gray-scale data is available, any local algorithm for Minkowski tensors using black-and-white voxels will be asymptotically biased, i.e., the estimator will not converge to the true value in the limit of infinite resolution [77].
@@
An algorithmic solution to the problem was recently proposed in [32], which had been inspired by [16, 56]. A more general framework, admitting to work with general distance-like functions (such as power distances and, correspondingly, power cells) to improve the robustness of estimation procedures against outliers, while preserving the resilience with respect to Hausdorff noise, was explored in [22]; for a discussion of and references to related methods, see [22]. The key idea of the algorithmic solution in [32] is to construct Voronoi cells that can encode global information and thus guarantee asymptotically bias-free estimators of all Minkowski tensors for sets of positive reach. This requirement of a positive reach, however, excludes many interesting applications, like rough surfaces and composite materials. More importantly, the proposed algorithm suffered from numerical instabilities since it relies on a matrix inversion, which is slow and sensitive to rounding errors, especially for ill‑conditioned matrices.
@@
Here, we substantially advance the robustness of the algorithm by replacing the matrix inversion by a least squares fit. The latter avoids numerical instabilities and allows for more data from parallel sets than needed, which further enhances robustness against statistical and systematic errors. Furthermore, we increase its efficiency by using an unbiased estimator of the Voronoi tensors (also called Voronoi volume integrals). This revised algorithm, here called Voronoi-LSQ algorithm, is also asymptotically unbiased for sets of positive reach, as we rigorously prove; see Remark 4.13. We demonstrate its robustness and accuracy for exemplary cases of geometric shapes, stochastic models, and real data.
@@
Even for a set without positive reach, our algorithm achieves surprisingly good results with relative deviations of less than 3% in the eigenvalues of the interfacial tensors, see Table 6, at least for the simple example of a rectangular shell. Since such precision is not theoretically substantiated, we also introduce an alternative method, here called Voronoi-FD algorithm, to estimate interfacial tensors, for which we are able to prove that its corresponding estimator is asymptotically unbiased for a large class of sets, including parallel sets of compact sets or finite unions of compact sets with positive reach (see Theorem 4.8).
@@
Together with the paper, we publish an open-source python package [60] including both methods. It is the first implementation of Minkowski tensors that we know of that can be applied to any dimension. It can also be applied to any discrete representation of the set, i.e., not only for pixelated images but representations using non-cubic lattice or even random point patterns. The last generalization is especially helpful in high dimensions, where, similarly to numerical integration, lattice representations may converge too slowly.
In Section 2, we provide an introduction to integral geometry and specifically to Minkowski tensors. In Section 3, we first further explain and illustrate the main theoretical foundations from geometric measure theory and then establish two results required for the algorithms presented in Section 4. Details on the algorithm, its theoretical foundation, implementation, and suitable choices of parameters are discussed in Sections 4.14.5.
@@
We demonstrate the reliability of our methods for simple geometric test cases in Sections 5.1 and 5.2 (see also Table 2). For a stochastic test case, we apply our estimator to a classic example of random convex sets, isotropic random polytopes (see Section 6), both in theory (see Section 6.1) and simulations (see Section 6.2).
@@
In Section 7, we turn to experimental data. First, we analyze metallic grains in a polycrystalline nickel-based superalloy using data from [74]. The Minkowski tensors characterize the cells as being distinctly more anisotropic with respect to curvature than the surface area (see Section 7.1). Then, we determine the interfacial tensors of nanorough surfaces from [73] (see Section 7.2).
@@ 2 Minkowski tensors in integral geometry
In this section, we provide a brief introduction to Minkowski tensors (or tensor valuations) and summarize their basic properties. For basic concepts from convex and integral geometry not introduced here, we refer to [35, 69]. Let ${\mathcal{K}}^{d}$ denote the set of all compact convex subsets (convex bodies) of ${\mathbb{R}}^{d}$. We denote by $\langle\cdot\,$$\cdot\rangle$ and $\|\cdot\|$ a Euclidean scalar product and the induced norm. We write $B^{d}:=\{x\in{\mathbb{R}}^{d}:\|x\|\leq 1\}$ for the Euclidean unit ball centered at the origin $o$ and $\mathbb{S}^{d-1}=\partial B^{d}=\{x\in{\mathbb{R}}^{d}:\|x\|=1\}$ for its boundary. We set $B^{d}(x,r):=x+rB^{d}$ for $x\in{\mathbb{R}}^{d}$ and $r\geq 0$. The volume of $B^{d}$ is denoted by $\kappa_{d}=\pi^{d/2}/\Gamma(1+d/2)$ and $\omega_{d}=d\kappa_{d}$ is its surface area, i.e., the (d1)(d-1)-dimensional volume of $\mathbb{S}^{d-1}$. The intrinsic volumes $V_{i}:{\mathcal{K}}^{d}\to[0,\infty)$, $i\in\{0,\ldots,d\}$, are a collection of $d+1$ basic functionals on ${\mathcal{K}}^{d}$ that are distinguished by their properties:
  • They form a basis of the vector space of continuous, isometry invariant, additive, functions on ${\mathcal{K}}^{d}$; $V_{i}$ is positively homogeneous of degree $i$ [35, Theorem 4.20].
  • They arise as coefficient functionals of a Steiner formula [35, Theorem 3.10].
  • They satisfy various integral geometric formulas such as Crofton formulas, projection formulas and (intersectional as well as additive) kinematic formulas [35, Chap. 5].
  • They can be additively extended to polyconvex sets (finite unions of compact, convex sets) and play a key role in the classical theory of geometric valuations (see [68, Chap. 1] and [35, Chap. 4.5]).
In particular, $V_{d}$ is the volume functional, $2V_{d-1}$ is the surface area, $V_{1}$ is proportional to the mean width functional and $V_{0}$ is the Euler characteristic (see [35, Section 3.3]). The intrinsic volumes have also been called Minkowski functionals (or quermassintegrals, though with a different normalization and notation). Due to their properties, the Minkowski functionals are useful descriptors that can be applied for the analysis of complex spatial structure.
Since the Minkowski functionals are isometry invariant, their usefulness in characterizing anisotropic features of objects under investigation is limited. The Minkowski tensors (see [30, Chap. 2]), which we introduce next, provide a more general set of tensor-valued geometric descriptors that are sensitive to position and orientation of geometric objects in Euclidean space (see Fig. 1 for an illustration). For detailed information on the determination and reconstruction of (classes of) convex bodies from volume or surface tensors we refer to [48, 46, 47] and to [33, Chap. 7]. We start by providing a brief introduction to symmetric tensors and tensor-valued functionals as needed for the present purpose.
@@
where $m_{k}:=|\{\ell\in\{1,\ldots,r\}:i_{\ell}=k\}|$, $k=1,\ldots,d$, counts how often the number $k$ appears among the indices $i_{1},\dots,i_{r}$. In particular, for $d=r=2$, we have $T(e_{1},e_{1})=t_{11}$, $T(e_{2},e_{2})=t_{22}$ and $T(e_{1},e_{2})=T(e_{2},e_{1})=\frac{1}{2}t_{12}$, hence $T=T(e_{1},e_{1})e_{1}^{2}+2T(e_{1},e_{2})e_{1}e_{2}+T(e_{2},e_{2})e_{2}^{2}$ with $2T(e_{1},e_{2})e_{1}e_{2}=T(e_{1},e_{2})e_{1}e_{2}+T(e_{2},e_{1})e_{2}e_{1}$.
@@
In order to introduce the basic Minkowski tensors of (nonempty) convex bodies, we use the support measures which can be considered as local versions of the intrinsic volumes. For a convex body $K\subset{\mathbb{R}}^{d}$, the support measure $\Lambda_{j}(K,\cdot)$, for $j\in\{0,\ldots,d-1\}$, is a Borel measure on ${\mathbb{R}}^{d}\times\mathbb{S}^{d-1}$ that arises as a coefficient in a local Steiner formula (compare (3) for a much more general relation) and satisfies $\Lambda_{j}(K,{\mathbb{R}}^{d}\times\mathbb{S}^{d-1})=V_{j}(K)$.
For instance, if $P\subset{\mathbb{R}}^{d}$ is a polytope, $\mathcal{F}_{j}(P)$ is the finite set of $j$-dimensional faces of $P$, and $N(P,F)$ is the (dj)(d-j)-dimensional normal cone of $P$ at $F\in\mathcal{F}_{j}(P)$, then
@@
$\Lambda_{j}(P,\cdot)=\frac{1}{\omega_{d-j}}\int_{F}\int_{N(P,F)\cap\mathbb{S}^{d-1}}\mathds{1}\{(x,u)\in\cdot\}\,\mathcal{H}^{d-1-j}(\mathrm{d}u)\,\mathcal{H}^{j}(\mathrm{d}x),$ (2.4)
where $\mathcal{H}^{k}$ denotes the $k$-dimensional Hausdorff measure, for $k\in\{0,1,\ldots,d\}$, on the appropriate domain, respectively (see Fig. 2 for an illustration of faces and normal cones of a polytope). For a face $F$ of a polytope $P$, the normal cone $N(P,F)$ is defined as the set of all $u\in{\mathbb{R}}^{d}$ such that $\langle z,u\rangle\leq\langle x,u\rangle$ for all $z\in P$, where $x$ is an arbitrary point in the relative interior of $F$ (see [69], Section 2.2). An alternative description is provided in [35] (Solution for Exercise 3.3.9). The pairs $(x,u)\in F\times N(P,F)$ are called support elements of $P$. Since support measures are concentrated on the support elements of the underlying set (also in the case of general sets), the terminology is justified.
Figure 2: Cube $P$ with normal cones $N(P,F_{i})$ at faces $F_{i}\in\mathcal{F}_{i}(P)$ for $i=0,1,2$.
Another basic description of the support measures is available for convex bodies (or more general domains) $K\subset{\mathbb{R}}^{d}$ with nonempty interior whose boundary $\partial K$ is of class $C^{2}$. If $\nu(K,x)$ denotes the unique exterior unit normal vector of $K$ at $x\in\partial K$, then
$\Lambda_{j}(K,\cdot)=\frac{1}{\omega_{d-j}}\int_{\partial K}\mathds{1}\{(x,\nu(K,x))\in\cdot\}\sum_{|I|=d-1-j}\prod_{i\in I}k_{i}(K,x)\,\mathcal{H}^{d-1}(\mathrm{d}{x}),$ (2.5)
where $k_{1}(K,x),\ldots,k_{d-1}(K,x)\geq 0$ are the principal curvatures of $K$ at $x\in\partial K$ and the summation extends over all subsets $I\subseteq\{1,\ldots,d-1\}$ of cardinality |I|=d1j|I|=d-1-j (if j=d1j=d-1 the empty product is interpreted as $1$).
For basic properties of the support measures, their connection to a local Steiner formula and a representation of $\Lambda_{j}(K,\cdot)$ for general convex bodies, which produces (2.4) and (2.5) as special cases, we refer to [69, (2.78) and (4.28)] and Section 3. The maps $K\mapsto\Lambda_{j}(K,\cdot)$ are measure-valued, additive (valuations), weakly continuous and for Borel sets $\alpha\subseteq{\mathbb{R}}^{d}$ and $\beta\subseteq\mathbb{S}^{d-1}$ they satisfy the covariance condition $\Lambda_{j}(gK,g\alpha\times g_{0}\beta)=\Lambda_{j}(K,\alpha\times\beta)$, where $g:{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$ is a rigid motion and $g_{0}$ is its rotational part (that is, $g(x)=g_{0}(x)+t$ for $x\in{\mathbb{R}}^{d}$, where $t\in{\mathbb{R}}^{d}$ is a translation vector and $g_{0}$ is a rotation). More generally, the support measures have been defined (as signed measures) for the more general class of sets of positive reach (see Section 3 for a definition and [63, Chap. 4] for a detailed introduction) and for finite unions of compact sets with positive reach all of whose finite intersections have again positive reach (see [63, Chap. 5] and the literature cited there); the corresponding class of sets in ${\mathbb{R}}^{d}$ is denoted by $\mathcal{U}^{d}$. In fact, extensions of these measures to arbitrary closed subsets of ${\mathbb{R}}^{d}$ have been studied in [28, 29] and will be considered in Section 3.
If $K$ is a convex body or a compact set with positive reach and $r,s\in{\mathbb{N}}_{0}$, then the basic Minkowski tensors of $K$ are defined by
$\Phi_{k}^{r,s}(K):=\frac{1}{r!s!}\frac{\omega_{d-k}}{\omega_{d-k+s}}\int_{{\mathbb{R}}^{d}\times\mathbb{S}^{d-1}}x^{r}u^{s}\,\Lambda_{k}(K,\mathrm{d}(x,u)),$ (2.6)
for $k\in\{0,\dots,d-1\}$,
@@
Support measures and thus Minkowski tensors can be defined not only for convex bodies or compact sets with positive reach. By additivity their domain can be extended to the class of polyconvex sets (finite unions of convex bodies). Moreover they can be defined also for sets from the class $\mathcal{U}^{d}$, since such an extension is possible for the support measures. Furthermore, Minkowski tensors can be introduced for even more general classes of sets provided that the support measures are defined in a natural way and the integrals exist (see Section 3 and the references given there). An explicit description of the Minkowski tensors of polytopes is based on (2.4) and used in Section 5.1, in Section 5.2 we evaluate formulas for nonconvex sets with smooth boundaries (compare (2.5)). In the case of sets with positive reach or for even more general classes of compact sets $K$ (such as arbitrary finite unions of sets with positive reach), the support measures $\Lambda_{j}(K,\cdot)$ are consistently replaced by the reach measures $\mu_{j}(K;\cdot)$ in definition (2.6); the relationship between these measures will be discussed in Section 3.
It is a straightforward consequence of the properties of the support measures of convex bodies that the Minkowski tensors of convex bodies are continuous, isometry (rigid motion) covariant (for a detailed definition, especially of the underlying notion of polynomial behavior with respect to translations, we refer to [30, Section 2.2]) and additive tensor functionals (valuations) on the space of convex bodies in ${\mathbb{R}}^{d}$. By a fundamental result due to Alesker [3, 4], the vector space $T_{p}$ of all mappings $\Gamma:\mathcal{K}^{d}\to\mathbb{T}^{p}$ having these properties is spanned by the tensor valuations $Q^{m}\Phi_{k}^{r,s}$, where $m,r,s\in{\mathbb{N}}_{0}$ satisfy $2m+r+s=p$, where $k\in\{0,\dots,d\}$, and where $s=0$ if $k=d$. In contrast to the real-valued and vector-valued case, these tensor valuations are no longer linearly independent. A study of linear dependencies was initiated by McMullen and completed by Hug, Schneider, Schuster (for precise references and detailed statements, see, e.g., [30, Theorems 2.6 and 2.7]). These general investigations imply that if $T_{2,k}$ denotes the subspace of all maps in $T_{2}$ that are homogeneous of degree $k$, then $T_{2}=T_{2,0}\oplus\cdots\oplus T_{2,d+2}$ (direct sum decomposition), $\dim T_{2}=3d+1$, and a basis of $T_{2,k}$ is displayed in Table 1. Denoting by $T_{2}^{\ast}$ the subspace of translation invariant maps in $T_{2}$ and by $T_{2,k}^{\ast}$ the subspace of translation invariant maps in $T_{2,k}$, we have $T_{2}^{\ast}=T_{2,0}^{\ast}\oplus\cdots\oplus T_{2,d}^{\ast}$, $\dim T_{2}^{\ast}=2d$, and a basis of $T_{2,k}^{\ast}$ is displayed in Table 1. Moreover, as a special consequence of McMullen’s linear dependencies, we obtain
$\Phi_{0}^{0,2s}=\chi\cdot(4\pi)^{-s}(s!)^{-1}Q^{s},$ (2.8)
for $s\in{\mathbb{N}}_{0}$, as indicated in Figure 1(c).